# Get a Grip: Reconstructing Hand-Object Stable Grasps in Egocentric Videos

### 25 Dec 2023 v1

@@ Abstract
We address in-the-wild hand-object reconstruction for a known object category in egocentric videos, focusing on temporal periods of stable grasps. We propose the task of Hand-Object Stable Grasp Reconstruction (HO-SGR), the joint reconstruction of frames during which the hand is stably holding the object. We thus can constrain the object motion relative to the hand, effectively regularising the reconstruction and improving performance.
By analysing the 3D ARCTIC dataset, we identify temporal periods where the contact area between the hand and object vertices remain stable. We showcase that objects within stable grasps move within a single degree of freedom (1 DoF). We thus propose a method for jointly optimising all frames within a stable grasp by minimising the object’s rotation to that within a latent 1 DoF.
We then extend this knowledge to in-the-wild egocentric videos by labelling 2.4K clips of stable grasps from the EPIC-KITCHENS dataset. Our proposed EPIC-Grasps dataset includes 390 object instances of 9 categories, featuring stable grasps from videos of daily interactions in 141 environments. Our method achieves significantly better HO-SGR, both qualitatively and by computing the stable grasp area and 2D projection labels of mask overlaps.
@@ 1 Introduction
Accurately reconstructing three-dimensional hands along with a grasped or manipulated object is key to unlocking many perception problems, including fine-grained understanding of interactions, but also potential applications including augmented reality, imitation learning and human-machine interactions.
Prior approaches have laid the foundation for the task, though focusing mostly on 3D input [41, 2, 53], which limits its application to the vast amount of video footage. Approaches with 2D input have relied on 3D ground-truth for supervision [51, 7, 19, 23, 50], again restricting their scope to curated setups where multi-camera sensing and 3D ground truth can be obtained. Early insights [19, 5] demonstrate that evaluation on in-the-wild egocentric footage remains significantly challenging.
In this work, we simplify the task of reconstruction in-the-wild by focusing on temporal periods of stable grasps. Our contribution is composed of four components. First, we propose the task of Hand-Object Stable Grasp Reconstruction (HO-SGR) which jointly optimises the reconstructions across all frames within one stable grasp. Second, using the ARCTIC dataset [13], we showcase that objects move within one degree of freedom (DoF) once within the hand’s stable grasp. Third, we accordingly propose a method that jointly reconstructs the hands and objects by minimising the object’s motion to 1DoF across frames. We demonstrate our method outperforms baselines and alternative assumptions of object movement using 3D ground truth from the stable grasps within the egocentric views of ARCTIC. Fourth, we focus on our objective of reconstruction for in-the-wild egocentric videos. We label a sizable dataset of 2.4K stable grasps clips from egocentric videos. Our EPIC-Grasps dataset is the first for hand-object reconstruction collected from unscripted activities, with individuals grasping 390 different objects by both hands. Similar to previous works [17, 6, 15, 5], we restrict our evaluation to known category CAD models. Our dataset comes with pseudo-ground truth in the form of 2D segmentation masks available from [12], allowing to measure the 3D reconstruction’s projection relative to this 2D groundtruth.
@@
We demonstrate a sample from our EPIC-Grasps dataset along with reconstructions from our method in Fig LABEL:fig:teaser.
@@ 2 Related Works
For comparison of hand object reconstruction datasets, see Sec 5. We review related works in object pose estimation, hand-object reconstruction and grasping.
Object Pose Estimation The task of object pose estimation serves as a foundation for the broader task of hand-object reconstruction. Object pose estimation involves estimating the 6D poses of objects from given CAD models. Labbe and Michaud [27] and Wang et al. [44] proposed instance-level methods for object pose estimation. However, the NOCS method proposed by Wang et al. [45] assumes that no exact CAD models are available and requires an additional size estimation step. To address this limitation, Liu et al. [29] have proposed a method that estimates the 3-DoF sizes of objects, resulting in a 9-DoF pose estimation method.
In this work, we model the object reconstruction as estimating 6D poses with a scalar size, as in [19, 45].
3D Hand-Object-Reconstruction 3D hand object reconstruction methods can be grouped into two categories.
The first category, known-CAD methods [19, 5, 33], assumes that object CAD models are given and fits 3D shapes into 2D observations. To address the limited availability of 3D annotations, several methods [18, 28, 48] have been proposed. Other methods [46, 43, 50, 1] explore hand-object interactions through different modeling approaches. HOMan [19] is a generalisation of single-frame methods that incorporates temporal knowledge through the smoothness of mesh vertices over time and the propagation of object pose initialization using the previous frame’s pose.
@@
The second category, CAD-agnostic methods, aims to estimate the hand and object poses without using explicit CAD models. An early attempt [42] models the object as a 3D bounding box and estimates its 6D pose. More recent methods, such as Diffusion-guided [52], ObMan [20] and implicit shape methods [51, 23, 7], use neural networks to encode the underlying object shape. However, recent methods like Huang et al. [22] require careful scanning of a firmly held object, making them impractical for use in in-the-wild videos.
Our dataset and method belong to the first category, which is a simplified assumption so we focus on the challenges of in-the-wild reconstruction.
3D Human Grasp Synthesis Generating plausible 3D human grasps for objects has been an active area of research.
Figure 1: We compare within/outside grasps, analysing object in-contact area (left), finger pose (middle) and relative object rotation w.r.t the hand (right), normalising all stable grasp duration for direct comparison (0 to 1 marked with red dotted lines). We showcase our manual annotations (top) and automatic (bottom).Objects do rotate within the stable grasp (blue) however mostly within 1 DoF (rotation residual in green).
GraspingField [23] characterises each point in a 3D space as the signed distances to the surfaces of the hand and the object, and learns a generative model to synthesize human grasps. ContactOpt [16] approximates deformable soft tissue in the hand by introducing a differentiable contact model. TOCH [56] learns to refine incorrect 3D hand-object interaction sequences using a correspondence-based prior. Zimeng et al. [55] utilises the physical stability to optimise single-image contact reconstruction.
However, as human grasp synthesis focuses on generating realistic grasps without incorporating input visual information, we do not explore these in this work.
3 What is a “stable grasp”?
The term “stable grasp” has been previously used in a number of prior works [4, 9, 14]. The term intuitively refers to a hand fully supporting the weight of the object, stopping it from slipping or sliding. During interactions, one hand holds an object in a “stable grasp” to either move it from one position to another, or for the second hand to manipulate the object. For example, if one is opening a bottle, then one hand is in a stable grasp while the other hand can be rotating the cap and is thus not in a stable grasp respect to the bottle.
Existing works on grasp type classification implicitly assume the given grasps are stable (or static) [14, 3, 35], without defining what a stable grasp is. Our first contribution is to explore what defines a stable grasp from a video sequence where the hand approaches an object, grasps it, firming the grip into a stable grasp and eventually releases the object. To do so, we leverage the recently introduced ARCTIC dataset [13]. The data offers a unique opportunity to study hand-object reconstruction as it includes 3D ground truth thanks to being captured in a MoCap system with 54 cameras and markers on both the hand and the objects.
For our exploration, we first analyse the footage and manually annotate 40 temporal extents of what we deem are stable grasps by visual inspection. Figure 1a shows that the stable grasp region corresponds to a consistent in-contact area during the temporal extent of the stable grasp—the same set of object vertices remain in contact with the hand during the grasp but change rapidly outside the grasp. We also demonstrate that both the finger tip poses and the object relative orientations do change during the stable grasps (average MPJPE [13] 4.6mm vs average rotation angle ${}^{\circ}$ for objects). However, these angles are significantly higher outside the stable grasp region (average MPJPE 11.8mm for fingers vs average rotation ${}^{\circ}$ for objects).
Figure 2: Sample stable grasp sequence from ARCTIC. Top: Dotted lines for start-end boundary. Bottom: stable contact area and object rotation along 1 DoF (purple).
Amongst the three metrics, we focus on the stable contact area as it gives the biggest change within/outside grasp. We thus define stable grasp as the temporal segment during which the object in-contact area has high intra-frame IOU
$\left[l^{*},r^{*}\right]=\underset{l,r}{\mathrm{argmax}}\left\{r-l\ :\mathrm{IOU}(S_{i},S_{j})>0.5\;\;\forall l\leq i<j\leq r\right\}$
where $S_{i}$ is the object vertices that are in-contact with the hand in frame $i$. That is the surface in-contact area should remain within an IOU of 0.5 to any other frames during the stable grasp. We use this definition to automatically find all stable grasp temporal boundaries in the ARCTIC dataset.
We accordingly locate 1303 stable grasp clips which we denote these as ARCTIC-Grasps dataset. In Fig 1b we compare this automatic set on the same metrics, showcasing consistent results. On average, 90% of frames have at least 74% IOU between any two frames in the stable grasp.
Object rotation within a stable grasp While object rotations within the stable grasp are consistently low (Fig 1 right), hands are so versatile that objects indeed do rotate, see Fig 2 for example. We observe that these rotations can be successfully approximated by rotations along a single degree-of-freedom (1 DoF). In Fig 1, we show the average rotation versus the rotation off a 1-DoF, calculated as the residual angle off this axis. This residual error is marginal within the stable grasp. In Sec 4.2 we elaborate on how we utilise this knowledge to optimise for hand object reconstructions within the stable grasp boundaries.
From 3D Datasets to Egocentric Videos. As noted in Sec 2 and our introduction, our objective is to perform HO-SGR in-the-wild without any 3D ground truth to assess the reconstruction. The absence of 3D ground truth has been an obstacle for evaluating performance in-the-wild, limiting it to qualitative results solely [52, 19]. By focusing on stable grasps, we offer a way forward to evaluation without 3D ground truth. We use the stable contact area as a measure of assessing successful reconstructions over the duration of the stable grasp. We then use the 1 DoF knowledge in HO-SGR, by minimising the rotation of the object off a latent axis of rotation, within which the object may still rotate. We explain our method next.
4 Method
Figure 3: Pipeline overview. We show 3 frames within a stable grasp. The yellow mesh color indicates 3D shapes that are being optimised; the blue mesh color indicates the underlying meshes are frozen. During the object relative pose optimisation stage, the object relative pose is optimised with 1-DoF constraint. During the hand-object refinement stage, the hand mesh and the object mesh are combined, and the ground-truth masks are merged.
4.1 Problem Statement and Notations
@@
Given a sequence of images of hand-object interactions, annotated with hand-side, known object category, start-end times of stable grasp and 2D segmentation of the hand and object, we aim to reconstruct consistent grasping, in every sampled frame $n$ of the sequence of length $N$, by producing a per-frame pair of 3D meshes of the hand and the object w.r.t the camera. We refer to this task as Hand-Object Stable Grasp Reconstruction (HO-SGR).
To represent the hand mesh, we use the MANO [36] model, which takes as input a finger articulation vector $\theta\in R^{45}$ and an identity-specific vector $\beta\in\mathbb{R}^{10}$, and outputs the hand mesh with vertices $V_{h}\in\mathbb{R}^{778\times 3}$ in the hand coordinate system. We use FrankMocap [37] to obtain MANO finger articulations $\theta^{n}$ from all the frames and we assume the mean hand shape, i.e. $\beta^{n}=0$. We thus have per-frame hand vertices $V_{h}^{n}=\mathrm{MANO}(\theta^{n},\beta^{n})$. The hand-to-camera (h2c) pose $T_{h2c}^{n}\in SE(3)$ [32], which is defined as the hand wrist orientation and position, are produce by FrankMocap by default and are used to transform $V_{h}^{n}$ to $V_{h:c}^{n}$ in the camera coordinate system for each frame.
@@
For the object mesh, we use $V_{o}\in\mathbb{R}^{|V_{o}|\times 3}$ to denote the known object vertices in the object coordinate system. We estimate the object-to-hand ($o2h$) poses $T_{o2h}^{n}\in SE(3)$ and the scale $s\in R$, which transform object vertices to $V_{o:h}^{n}$ in the hand coordinate system for each frame. We then use the hand-to-camera (h2c) pose $T_{h2c}^{n}$ to transform $V_{o:h}^{n}$ to $V_{o:c}^{n}$ in the camera coordinate system for each frame.
@@
4.2 Reconstructing Objects in a Stable Grasp
The stable grasp segments are defined based on the in-contact surface area, yet it is non-trivial to directly incorporate the Stable Contact Area constraint into optimisation, as the contact area is not known in advance.
Instead, as shown in Sec 3, we constrain the reconstruction so the in-hand object motion is minimised to a single degree-of-freedom, i.e. the object rotates along a fixed latent axis. We represent the 6D object-to-hand transform as a composition of transformation functions
$T_{o2h}^{n}=P_{o2h}\circ P^{-1}_{o2\phi}\circ R_{\phi}^{n}\circ P_{o2\phi}$ (3)
i.e. a per-frame rotation $R_{\phi}^{n}$ is first applied to the object along the latent axis, defined by $P_{o2\phi}$, then the common object-to-hand transform $P_{o2h}$ is applied. Furthermore, this does not restrict the ability of the hand to move relative to the camera or the wrist to move freely around the joint.
By back-propagating the error through $V_{o:h}^{n}$, the common parameters across all frames ($P_{o2h}$, $P_{o2\phi}$, $s$) as well as the individual frame parameters $\{R_{\phi}^{n}\}$ will be optimised jointly.
4.3 Initial Hand Mesh Refinement
Before optimising the object pose, we perform two steps of refinements for the FrankMocap hand output: temporal smoothness and 2D hand mask. We describe these next.
We define a general temporal smoothness function:
$F_{s}(\{\boldsymbol{x}^{n}\})=\sum_{n=2}^{N-1}|\boldsymbol{x}^{n}-\frac{\boldsymbol{x}^{n-1}+\boldsymbol{x}^{n+1}}{2}|^{2}_{2}$ (4)
We first optimise the finger articulation vector $\theta^{n}$ with the above temporal smoothness objective:
$E(\{\theta^{n}\})=F_{s}(\{\theta^{n}\})$ (5)
Then, we use the projected 2D hand mask to optimise the hand pose at every frame. We optimise the hand-to-camera pose $T_{h2c}^{n}=[R_{h}^{n}|t_{h}^{n}]$ for the hand mask pixel difference with ground-truth as well as temporal smoothness of poses:
$E(\{R_{h}^{n}\},\{t_{h}^{n}\})$ $=F_{s}(\{R_{h}^{n}\})+F_{s}(\{t_{h}^{n}\})+\sum_{n=1}^{N}E_{mask}^{n}$ (6)
$E_{mask}^{n}$ $=|\mathcal{C}_{h}^{n}\otimes(\mathcal{M}_{h}^{n}-\Pi(V_{h:c}^{n}))|^{2}_{2}$ (7)
where $\Pi(\cdot)$ is the differentiable projection function [24, 34] and $\mathcal{M}_{h}^{n}$ is the hand mask. $\mathcal{C}_{h}^{n}$ is the occlusion-aware mask as in [19, 54] which only computes the error within regions of the hand that are not occluded by the grasped object $\mathcal{C}_{h}^{n}$, set to 1 for the hand and background, and 0 for the object.
4.4 Optimising Object Relative Pose
After refining the hand wrist pose $T_{h2c}^{n}$ and finger articulation $\theta^{n}$ in Sec 4.3, we keep these fixed, and optimise the object relative pose parameters ($P_{o2h}$, $P_{o2\phi}$, $\{R_{\phi}^{n}\}$) and scale $s$ by minimising the following objective function:

### 07 Apr 2024 v2

# https://zhifanzhu.github.io/getagrip

@@ Abstract
We propose the task of Hand-Object Stable Grasp Reconstruction (HO-SGR), the reconstruction of frames during which the hand is stably holding the object. We first develop the stable grasp definition based on the intuition that the in-contact area between the hand and object should remain stable. By analysing the 3D ARCTIC dataset, we identify stable grasp durations and showcase that objects in stable grasps move within a single degree of freedom (1-DoF). We thereby propose a method to jointly optimise all frames within a stable grasp, minimising object motions to a latent 1-DoF. Finally, we extend the knowledge to in-the-wild videos by labelling 2.4K clips of stable grasps. Our proposed EPIC-Grasps dataset includes 390 object instances of 9 categories, featuring stable grasps from videos of daily interactions in 141 environments. Without 3D ground truth, we use stable contact areas and 2D projection masks to assess the HO-SGR task in the wild. We evaluate relevant methods and our approach preserves significantly higher stable contact area, on both EPIC-Grasps and stable grasp sub-sequences from the ARCTIC dataset.
Keywords:
Hand-Object Reconstruction 3D Reconstruction from Video Egocentric Vision Stable Grasp Dataset
@@ 1 Introduction
Accurately reconstructing three-dimensional hands along with a grasped or manipulated object is key to unlocking many perception problems, including fine-grained understanding of interactions, but also potential applications including augmented reality, robotic imitation learning and human-machine interactions.
Prior approaches have laid the foundation for the task, though focusing mostly on 3D input [39, 2, 49], which limits its application to the vast amount of video footage. Approaches with 2D input have relied on 3D ground truth for supervision [47, 6, 18, 22, 46], again restricting their scope to curated setups where multi-camera sensing and 3D ground truth can be obtained. Early insights [18, 4, 31] demonstrate that evaluation on in-the-wild egocentric footage remains significantly challenging.
In this work, we focus on the task of reconstruction in-the-wild on temporal periods of stable grasps. Our contribution is composed of three components. First, we propose the task of Hand-Object Stable Grasp Reconstruction (HO-SGR) which jointly optimises the reconstructions across all frames within one stable grasp. We showcase that objects move within one degree of freedom (1-DoF), relative to the hand pose, throughout the stable grasp. Second, we accordingly propose a method that jointly reconstructs the hands and objects by minimising the object’s motion, relative to the hand, to 1-DoF around a latent rotation axis, throughout the frames. We demonstrate our method outperforms baselines and alternative assumptions of object movement using 3D ground truth from the stable grasps within the egocentric views of the ARCTIC dataset [13]. Third, We label a sizeable dataset of 2.4K stable grasps clips from egocentric videos. Our EPIC-Grasps dataset is the first for hand-object reconstruction collected from unscripted activities, with individuals grasping 390 different objects by both hands. Similar to previous works [17, 5, 15, 4, 31], we restrict our evaluation to known category CAD models. Our dataset comes with pseudo-ground truth in the form of 2D segmentation masks available from [11], allowing to measure the 3D reconstruction’s projection relative to this 2D ground truth.
@@
Figure 1: Two stable grasp sequences from EPIC-Grasps for a bottle (left) and bowl (right). We show sample frames (top) and reconstructions (bottom). Right: for each reconstruction, we show the rotated view, along with the latent 1-DoF axis.
We demonstrate a sample from our EPIC-Grasps dataset along with reconstructions from our method in Fig. 1.
@@ 2 Related Works
For comparison of hand object reconstruction datasets, see Sec. 4. We review related works in hand pose estimation and hand-object reconstruction.
3D Hand Pose Estimation. Estimating 3D hand pose from RGB images has been proposed for both free hands and hands in-interactions. FrankMocap [35] is commonly used CNN-based method and has been integrated into many hand-object reconstruction methods [18, 4, 31, 47, 48]. METRO [26] proposes to use a transformer on top of the CNN feature for regression. HaMeR [32] is a recent transformer-based method with improved training data scale and higher network capacity. A concurrent work, WildHands [33], addresses in-the-wild egocentric hand pose estimation by considering perspective distortion and trains on relevant in-the-wild data. In this work, we use HaMeR[32] due to its superior performance. We also show HaMeR leads to better stable grasp reconstruction than FrankMocap in ablations.
3D Hand-Object-Reconstruction. Methods are grouped into two categories.
The first category, known-CAD methods, assumes that object CAD models are given and fits 3D shapes into 2D observations. These can further be classified into learning-based [40, 27, 44, 42, 41, 46, 1] or optimisation-based [18, 4, 31] methods. Learning-based methods learn to jointly reconstruct hands and objects from seen object examples, whereas optimisation-based methods address the reconstruction by directly fitting to 2D signals. RHO [4] is the first optimisation based single-frame method. The optimisation-based methods [4, 18, 31] share the same pipeline where hand/object is first independently optimised, followed by joint optimisation with physical terms. In particular, HOMan [18] is a generalisation of the single-frame method that incorporates temporal knowledge through the smoothness of mesh vertices over time and the temporal propagation of object pose initialisation.
@@
The second category, CAD-agnostic methods, aims to estimate the hand and object poses without using explicit CAD models. Many CAD-agnostic methods [48, 19, 47, 22, 6] learn object shape priors from the data. The other line of methods [21, 16, 12] uses neural networks to fit the underlying object shape from multiple views. However, these methods typically require careful scanning of a held object, making them impractical for use in in-the-wild videos.
Our dataset and method belong to the first category, which is a simplified assumption needed for the challenges of in-the-wild reconstruction. Different from above optimisation-based methods, our approach examines the object’s relative motion and proposes to constrain the relative motion during stable grasp.
3 Stable Grasp Reconstruction: Problem and Method
We first define the standard task of hand-object reconstruction, then detail the adjustments we do to focus on stable grasps.
@@ 3 Stable Grasp Reconstruction: Problem and Method
The general task of Hand-Object Reconstruction is defined as: given a sequence of $N$ images of hand-object interactions, annotated with hand-side (left v.s. right) and a known object category, the task aims to produce, for every frame $n$, a pair of 3D meshes of the hand and the object w.r.t. the camera.
Different from the general task, we focus on a special temporal segment of object interactions – the stable grasp —which we will define in Sec 3.2. Given a start-end segment of a stable grasp, we aim to produce consistent hand object reconstructions across all frames within the stable grasp. We refer to this special-case task as Hand-Object Stable Grasp Reconstruction (HO-SGR).
3.1 Background and Notations
[…]
@@ 3.1 Background and Notations
For the object mesh, we use $V_{o}\in\mathbb{R}^{|V_{o}|\times 3}$ to denote the known object vertices in the object coordinate system. Different from [18, 31, 13], which estimates the object w.r.t. camera, we estimate the object-to-hand ($o2h$) poses $T_{o2h}^{n}\in SE(3)$ and the scalar scale $s\in\mathbb{R}$, which transforms object vertices to $V_{o:h}^{n}$ in the hand coordinate system for each frame. We then use the hand-to-camera (h2c) pose $T_{h2c}^{n}$ to transform $V_{o:h}^{n}$ to $V_{o:c}^{n}$ in the camera coordinate system for each frame.
@@
3.2 What is a stable grasp (SG)?
SG Definition. The term stable grasp has been previously used in human grasp analysis [3, 8, 14]. While definitions vary, they centre around the object being “held securely with one hand, irrespective of the hand orientation” [14]. Intuitively, this definition implies the hand maintains a stable contact (area) with the object – i.e. the same vertices of objects and hands are in contact for the duration of the grasp. For example, if a hand pours liquid out of a bottle, the hand indeed maintains contact with the same vertices in the bottle throughout manipulation, tilting to pour, then tilting back again upright before putting down. While the hand orientation and finger poses change during this interaction, the key here is the consistent stable contact area, which can be quantitatively computed from hand-object meshes. We visualise this example in Fig. 2.
Figure 2: Sample hand-object mesh sequence from ARCTIC. Contact areas (in shiny yellow) are similar within the stable grasp (blue background). In -0.16s the hand has no contact with the object.
Formally, for any pair of frames $i$ and $j$ within the stable grasp, we use $S_{i}$ and $S_{j}$ to denote the in-contact area on the object surface, and intersection-over-union $\mathrm{IOU}(S_{i},S_{j})$ between in-contact areas. Following above intuition, the duration of the stable grasp is defined as
$\left[l^{*},r^{*}\right]=\underset{l,r}{\mathrm{argmax}}\left(r-l\right)\quad\textrm{s.t.}\;\mathrm{IOU}(S_{i},S_{j})>\tau\;\;\forall l\leq i<j\leq r$ (2)
where $\tau$ specifies the minimum IOU threshold. The $argmax(r-l)$ implies the longest duration representing the stable grasp, from its initiation to conclusion.
SG Object Motion. It is critical to note that during the stable grasp, the hand still has dexterity to move or rotate an object relative to the hand pose; in other words, the object has non-static motions w.r.t. the hand. We visualise this distinction in Fig. 3 where we plot one manipulation sequence using two assumptions. The top assumes the object remains static relative to the hand pose – i.e. when the hand coordinate systems are aligned over the sequence, the object is perfectly aligned. This top example is actually a made-up one to explain the concept. The bottom example is a true sequence from the ARCTIC [13] dataset. We showcase that when the hand coordinate system is aligned over time, the object is visibly moving relative to the hand. The figure also shows that this motion is not completely free, but is restricted to rotation around a latent axis $\phi$. This finding about the object’s motion during the stable grasp is one of our paper’s contributions. We detail next the study we carried out, which resulted in this finding.
Figure 3: To study the object’s relative pose, we align the hand coordinate systems (right). Top: Made-up sequence with static relative pose – object is perfectly aligned (mixture of 3 colours). Bottom: True sequence showcasing object’s motion relative to the hand can be approximated as 1-DoF rotation around axis $\phi$, shown in purple.
SG Study. We perform a study on the 3D-ground-truth dataset ARCTIC [13] to analyse different quantitative measures within/outside the temporal extent of the stable grasp. We use the threshold $\tau=0.5$ (Eq 2) and automatically extract 1303 stable grasp sequences from a variety of objects and subjects throughout dataset (see appendix for details).
Figure 4: We compare within/outside grasps, analysing object in-contact area (left) and corresponding rotation errors of the static and 1-DoF rotation approximations (right), normalising all stable grasp duration for direct comparison (0 to 1 marked with blue background). While both the Static and 1-DoF assumptions result in low approximation error within stable grasps, the error of 1-DoF assumption is marginal (right).
We present the main finding in Fig 4. As anticipated, we show that the in-contact IOU, drops sharply outside the stable grasp – plotted in Fig 4 (left). That is the contact area remains stable only within the stable grasp’s temporal extent. We similarly compare other quantities (e.g. hand pose changes, finger-tip pose variations and object rotation) within/outside stable grasps (details in appendix). When we assess the relative object motion to the hand coordinate system within/outside the stable grasp, we note the findings first presented in Fig 3. Formally, we define a rotation axis $\phi$ around which the object can rotate – we detail how we find this axis later. If an object is only allowed to rotate around this axis, the motion is restricted from its free 6-DoF to a single rotation angle around this axis – we thus refer to this as a 1-DoF motion. The object pose w.r.t. the hand would then be described as
$T_{o2h}^{n}=\mathbf{rot}\left(\omega^{n},\phi\right)\circ T_{o2h}$ (3)
i.e., we first apply one global object-to-hand transform $T_{o2h}$ for all frames, followed by applying the per-frame rotation $\mathbf{rot}\left(\omega^{n},\phi\right)$ – the rotation of angle $\omega^{n}$ around the given rotation axis $\phi$. Here $\circ$ denotes composition of transformations.
Given the ground truth hand and object poses across a sequence, we can find the optimal rotation axis $\phi$ by minimising the error between the predicted object-to-hand transformation and the ground truth ones, that is:
$\underset{\phi,\{\omega^{n}\},T_{o2h}}{\mathrm{argmin}}\frac{1}{N}\sum_{n=1}^{N}\left\|T_{o2h}^{n}-T_{gt}^{n}\right\|^{2}$ (4)
where $T_{gt}^{n}$ is the ground truth object-to-hand transformation. In Fig. 4 (right) we plot this within/outside the temporal extent of stable grasp and compare that to the assumption that objects remain static (i.e. does not move relative to the hand). We calculate two rotation errors deviated from the ground truth – the first assumes the object remains static, and the second allows the object to move freely around $\phi$ and calculates the angle error off this axis. We show that the angle error when using the static assumption to be non-negligible (avg $10^{\circ}$). When we use the 1-DoF assumption, the error off the rotation axis $\phi$ is generally low (avg $3^{\circ}$).
In appendix, we further confirm these findings on the HOI4D [28] dataset. This finding on the object’s motion within the SG forms the base of our proposed optimisation for HO-SGR, which we present next.
3.3 Reconstructing Object Poses in a Stable Grasp
Figure 5: Our proposed reconstruction method. We show 3 frames within a stable grasp. HaMeR [32] produces the hand meshes (rendered in blue) from RGB, and we set the object-to-hand pose $T_{o2h}^{n}$ to the same $T_{o2h}$ initially. Then, during each iteration of the optimisation, the object’s relative pose is optimised to 1-DoF and projected back to individual frames. These are compared with ground truth segmentation (right), jointly optimise for all frames. We ignore mask computation in hand occluded region (grey in the right figure). The physical terms are omitted in this figure.
Recall that we aim to reconstruct the $N$ pairs of 3D meshes of the hand and the object during the stable grasps. We use HaMeR [32] to reconstruct the hand pose individually in every frame. HaMeR is a fully transformer-based architecture of high accuracy and robustness, trained on very large scale hand data. Given an RGB image, HaMeR takes as input the hand detection bounding box with a known hand side, and outputs the MANO finger articulations $\theta^{n}$ and the hand-to-camera pose $T_{h2c}^{n}$.
We then reconstruct the object relative pose with the render-and-compare approach, overviewed in Fig. 5. We propose to leverage our finding on the stable grasp, to optimise the object pose relative to the hand. It is non-trivial to directly incorporate the stable contact area into the optimisation, as the contact area is not known in advance. Instead, we use our finding on the object relative motion, and minimise this to a single degree-of-freedom (1-DoF), i.e. the object rotates along a fixed latent axis to-be-optimised. As shown in Sec. 3.2, the 1-DoF object relative pose $T_{o2h}^{n}$ at frame $n$ is given by per-frame rotation $\omega^{n}$, the rotation axis $\phi$ and the global base object-to-hand pose $T_{o2h}$ that takes the object into the hand.
Our main objective function is given by: